OpenStudy (anonymous):

Evaluate each polynomial for x = -1 and y = 3. 5x – y

OpenStudy (anonymous):

HELP!! Please! thanks! :)

OpenStudy (anonymous):

5(-1) - 3 = -8

OpenStudy (hoblos):

replace x by (-1) and y by (3) 5(-1) - (3) = -5-3 = -8

OpenStudy (anonymous):

5(-1)-3=-8

OpenStudy (anonymous):

Thanks! :)

OpenStudy (hoblos):

you are welcome :)